OpenStudy (anonymous):

scientific notation simplify 65x10^3

OpenStudy (anonymous):

65000

OpenStudy (ash2326):

you should have only one term to the left of decimal point so we get \[\large 65\times 10^3=6.5 \times 10\times 10^3=6.5\times 10^4\]

OpenStudy (ash2326):

if I have 10 its scientific form will be \[1\times 10^1\]

OpenStudy (anonymous):

so the exponent increase right

OpenStudy (ash2326):

yes but if you have \[0.1\] then its scientific notation will be \[1\times 10^{-1}\]